# Regularization in Data-driven Predictive Control: A Convex Relaxation Perspective

### v1

@@ 7 Conclusion
In this paper, we have analyzed the role of regularizations in direct and indirect data-driven control via a bi-level optimization framework. We prove that, after dropping some inner constraints, the bi-level optimization problem can be transformed into a single-level convex optimization problem. Moreover, regularizers developed from those single-level problems are exact penalty functions under certain conditions. Thus, using regularizers is a further convex relaxation with respect to the original bi-level problem (after dropping constraints). Moreover, we have proposed a new variant, called O-DDPC, which uses an iterative algorithm to obtain a novel data-driven control approach. Numerical simulations have demonstrated the superior performance of O-DDPC (38) with the more structured predictor. Future directions include investigating the closed-loop performance across various DDPC variants, and validating the proposed O-DDPC approach through real-world experiments.
@@ Appendix A Exact penalty for quadratic optimization
The proof for Theorem 1 is divided into two parts:
  1. 1.
    We transform (17) and (18) to two new optimization problems, where the optimal solutions $x_{1}^{*}$ remains the same.
  2. 2.
    We next establish an exact penalty such that new formulations have the same optimal solutions with a finite regularization parameter.
We first change the variables in (17) and (18) to facilitate the penalty analysis. The key idea is to construct a new decision variable $\bar{x}_{2}$ such that $\bar{x}_{2}=Dx_{2}$ with a free decision variable $x_{3}$. We consider two problems
$\min_{x_{1}\in\mathcal{X},\bar{x}_{2}\in\mathcal{D},x_{3}}$ $x_{1}^{{\mathsf{T}}}Mx_{1}$ (40)
$\mathrm{subject~to}$ $A_{1}x_{1}+A_{2}D^{\dagger}\bar{x}_{2}+A_{2}D^{\bot}x_{3}=b,$
$\|\bar{x}_{2}\|_{p}=0,$
and its penalized version
$\min_{x_{1}\in\mathcal{X},\bar{x}_{2}\in\mathcal{D},x_{3}}$ $x_{1}^{{\mathsf{T}}}Mx_{1}+\lambda_{w}\|\bar{x}_{2}\|_{p}$ (41)
$\mathrm{subject~to}$ $A_{1}x_{1}+A_{2}D^{\dagger}\bar{x}_{2}+A_{2}D^{\bot}x_{3}=b,$
where columns of $D^{\bot}$ form a basis for the null space of $D$ and $\mathcal{D}:=\mathrm{Im}(D)$. We have the following equivalence among optimization problems (17), (18), (40) and (41)
Proposition 6.
Problem (17) and (40) share the same optimal solution $x_{1}^{*}$, and problem (18) and (41) also share the same optimal solution $x_{1}^{*}$ for any $\lambda_{w}\geq 0$.
@@
We first show the equivalence of (17) and (40):
  1. 1.
    (17) and (40) have the same feasible region: if $x_{1},x_{2}$ is feasible to (17), we can construct $\bar{x}_{2}$ and $x_{3}$ such that $x_{1},\bar{x}_{2},x_{3}$ is feasible to (40). On the other hand, given any feasible solution $x_{1},\bar{x}_{2},x_{3}$, there exists $x_{2}$ such that $x_{1},x_{2}$ is feasible to (17).
  2. 2.
    The (17) and (40) have the same cost function over their feasible region.
Let $x_{1},x_{2}$ be a feasible solution to (17). We construct $\bar{x}_{2}$ as $\bar{x}_{2}=Dx_{2}$ and $x_{3}$ such that $D^{\bot}x_{3}=x_{2}-D^{\dagger}\bar{x}_{2}=x_{2}-D^{\dagger}Dx_{2}.$ The existence of $x_{3}$ is guaranteed as $D^{\dagger}Dx_{2}$ represents the projection of $x_{2}$ onto the row space of $D$, so that the residual must be included in the range space of $D^{\bot}$. Thus, we have $\|\bar{x}_{2}\|_{p}=\|Dx_{2}\|_{p}=0$ and
$A_{1}x_{1}+A_{2}D^{\dagger}\bar{x}_{2}+A_{2}D^{\bot}x_{3}=A_{1}x_{1}+A_{2}x_{2}=b,$
which means $x_{1},\bar{x}_{2},x_{3}$ is a feasible solution of (40).
We then let $x_{1},\bar{x}_{2},x_{3}$ is be feasible solution of (40). We now let $x_{2}=D^{\dagger}\bar{x}_{2}+D^{\bot}x_{3}$, and write
$\|Dx_{2}\|_{p}=\|DD^{\dagger}\bar{x}_{2}+DD^{\bot}x_{3}\|_{p}=\|\bar{x}_{2}\|_{p}=0,$
$A_{1}x_{1}+A_{2}x_{2}=A_{1}x_{1}+A_{2}D^{\dagger}\bar{x}_{2}+A_{2}D^{\bot}x_{3}=b,$
This shows $x_{1},x_{2}$ is a feasible solution for (17). Similarly, we can prove the equivalence of (18) and (41).  ∎
We next prove that $\|\bar{x}_{2}\|_{p}$ is an exactly penalty function for (40). Consider a perturbed version of (40):
$\min_{x_{1}\in\mathcal{X},\bar{x}_{2}\in\mathcal{D},x_{3}}$ $x_{1}^{{\mathsf{T}}}Mx_{1}$ (42)
$\mathrm{subject~to}$ $A_{1}x_{1}+A_{2}T^{\dagger}\bar{x}_{2}+A_{2}T^{\bot}x_{3}=b,$
$\|\bar{x}_{2}\|_{p}=\epsilon.$
We will use (42) to establish the following proposition.
Proposition 7.
Suppose $\mathcal{X}$ is a convex set, $M$ is positive semidefinite and $x_{1}^{{\mathsf{T}}}Mx_{1}$ is locally Lipchitz continuous with Lipchitz constant $L$ with respect to $\|\cdot\|_{p}$ over $\mathcal{X}$. If there exists an optimal solution $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ for (40) such that $x_{1}^{*}$ is an interior point of $\mathcal{X}$, there exists $\lambda_{w}^{*}>0$, such that (40) and (41) have same optimal solutions for all $\lambda_{w}>\lambda_{w}^{*}$.
Proof.
The key idea is to prove that the optimal solution $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ of (40) is a local minimum of (41) via the usage of (42). This optimal solution also becomes the global minimum since (41) is a convex problem. This is inspired by the notion of partial calmness [44].
  1. 1.
    Problem (42) permits a certain degree of “violation” of the constraint $\|\bar{x}_{2}\|_{p}=0$. We first show that, in a local neighborhood of $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$, such violations provide no benefit once incorporated into the objective function with an appropriate weighting parameter. This implies that $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ is a local optimal solution of (42).
  2. 2.
    Next, we establish that there exists a local region around $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ in which every feasible solution of (41) is also feasible for (42). Combining this fact with the result of Step 1, we conclude that $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ is a local optimal solution of (41). It is thus globally optimal due to the convexity of (41)
  3. 3.
    Finally, we demonstrate that all optimal solutions of (41) are also optimal solutions of (40).
Let $A_{3}:=[A_{1},A_{2}D^{\bot}]$ and $P_{2}:=\textnormal{col}(I,\mathbb{0})$. As $x_{1}^{*}\in\mathcal{X}^{\circ}$, there exists $\delta_{r}>0$ such that $\mathcal{B}_{\delta_{r}}(x_{1}^{*})\subseteq\mathcal{X}$. Let $\frac{\delta_{r}}{\sigma_{r}+\sigma_{p}}>\delta>0$ where $\sigma_{r}$ is the induced $p$-norm of $P_{2}A_{3}^{\dagger}A_{2}D^{\dagger}$ and $\sigma_{p}$ is the constant such that $\|v\|_{p}\leq\sigma_{p}\|v\|_{2}$ for any vector $v$ with the same dimension of $x_{1}$. For any $\epsilon\in[0,\delta)$, let $(x_{1},\bar{x}_{2},x_{3})\in\mathcal{B}_{\delta}(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ be a feasible solution for (42). We first show that there exists $\mu$, such that
$x_{1}^{{\mathsf{T}}}Mx_{1}+\mu\epsilon\geq x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}.$ (43)
We can represent $\textnormal{col}(x_{1},x_{3})$ as
$\textnormal{col}(x_{1},x_{3})=A_{3}^{\dagger}(b-A_{2}D^{\dagger}\bar{x}_{2})+v_{1},$
where $v_{1}\in\mathrm{Null}(A_{3})$. A feasible solution $(\tilde{x}_{1},\bar{x}_{2}^{*},\tilde{x}_{3})$ for (40) can be constructed as
$\textnormal{col}(\tilde{x}_{1},\tilde{x}_{3})=A_{3}^{\dagger}(b-A_{2}D^{\dagger}\bar{x}_{2}^{*})+v_{1}.$
We then demonstrate $\tilde{x}_{1}\in\mathcal{X}$ via deriving the difference between $x_{1}$ and $\tilde{x}_{1}$. We have
$\|x_{1}-\tilde{x}_{1}\|_{p}=\|P_{2}A_{3}^{\dagger}A_{2}D^{\dagger}\bar{x}_{2}\|_{p}\leq\sigma_{r}\epsilon\leq\sigma_{r}\delta.$
We then obtain
$\|x_{1}^{*}-\tilde{x}_{1}\|_{p}\leq\|x_{1}^{*}-x_{1}\|_{p}+\|x_{1}-\tilde{x}_{1}\|_{p}\leq(\sigma_{r}+\sigma_{p})\delta\leq\delta_{r},$
which implies $\tilde{x}_{1}\in\mathcal{B}_{\delta_{r}}(x^{*})\subseteq\mathcal{X}$. From the optimality condition of (40), we have
$x_{1}^{{\mathsf{T}}}Mx_{1}-x^{*{{\mathsf{T}}}}Mx^{*}\geq$ $x_{1}^{{\mathsf{T}}}Mx_{1}-\tilde{x}_{1}^{{\mathsf{T}}}M\tilde{x}_{1}$
$\geq$ $-L\|x_{1}-\tilde{x}_{1}\|_{p}\geq-L\sigma_{r}\epsilon,$
which means $x_{1}^{{\mathsf{T}}}Mx_{1}+L\sigma_{r}\epsilon\geq x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}.$ Thus, we can let $\mu\geq L\sigma_{r}$ so that the inequality (43) is satisfied.
We then prove that the optimal solution $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ of (40) is a local minima of (41). We can let $\lambda_{w}>\lambda_{w}^{*}\geq L\sigma$. Since $f(x):=\|x\|_{p}$ is a continuous function, there exists $\delta_{x}>0$ such that $\|\bar{x}_{2}\|_{p}<\delta$ for any $\bar{x}_{2}\in\mathcal{B}_{\delta_{x}}(\bar{x}_{2}^{*})$. Let $\delta^{*}:=\min(\delta_{x},\delta)$. For any feasible solution $(x_{1},\bar{x}_{2},x_{3})\in\mathcal{B}_{\delta^{*}}(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ of (41), there exists $\epsilon\in[0,\delta)$ such that it is feasible to (42). Thus, from the inequality (43), we have for any $(x_{1},\bar{x}_{2},x_{3})\in\mathcal{B}_{\delta^{*}}(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$,
$x_{1}^{{\mathsf{T}}}Mx_{1}+\lambda_{w}\|\bar{x}_{2}\|_{p}\geq x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*},$
which means $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ is a local minima (as well as the global minima) of (41).
It is now clear that all optimal solutions of (40) are optimal to (41). Next, suppose $(x_{1w}^{*},\bar{x}_{2w}^{*},x_{3w}^{*})$ is an optimal solution of (41) and recall that $\lambda_{w}>\lambda_{w}^{*}\geq L\sigma_{r}$. We have
$x_{1w}^{*{{\mathsf{T}}}}Mx_{1w}^{*}+\lambda_{w}\|\bar{x}_{2w}^{*}\|_{p}=x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}\leq x_{1w}^{*{{\mathsf{T}}}}Mx_{1w}^{*}+\lambda_{w}^{*}\|\bar{x}_{2w}^{*}\|_{p},$
which leads to $(\mu-\bar{\mu})\|\bar{x}_{2w}^{*}\|_{2}\leq 0$ and implies $\|\bar{x}_{2w}^{*}\|_{2}=0$. Thus, $(x_{1w}^{*},\bar{x}_{2w}^{*},x_{3w}^{*})$ is also feasible for (40) and is its optimal solution.  ∎
We combine Propositions 6 and 7 to establish Theorem 1.
Proof of Theorem 1: From Proposition 6, the optimal solution $x_{1}^{*}$ to (17) and (40) is the same. Furthermore, the assumption that $(x_{1}^{*},x_{2}^{*})$ is an optimal solution with $x_{1}^{*}\in\mathcal{X}^{\circ}$ for (17) implies there exists an optimal solution $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ and $x_{1}^{*}\in\mathcal{X}^{\circ}$ for (40). Thus, using Proposition 7, there exists $\lambda_{w}^{*}$ such that the optimal solution sets $x_{1}^{*}$ for (40) and (41) are equivalent for all $\lambda_{w}>\lambda_{w}^{*}$. That means those of (17) and (41) are the same with $\lambda_{w}^{*}$. Since we also have the optimal solutions of $x_{1}^{*}$ are the same for (18) and (41), this illustrates (17) and (18) have the same optimal solution set for $x_{1}^{*}$ with $\lambda_{w}^{*}$.
From Propositions 6 and 7, the optimal values of (17) and (18) are the same. As any optimal solutions $(x_{1}^{*},x_{2}^{*})$ for (17) are feasible for (18), substituting $(x_{1}^{*},x_{2}^{*})$ into the objective function of (18) we have
$x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}+\lambda_{w}\|Dx_{2}^{*}\|_{p}=x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*},$
which means $(x_{1}^{*},x_{2}^{*})$ is an optimal solution for (18). Then, suppose $(x_{1}^{*},x_{2}^{*})$ is an optimal solution for (18). Since $x_{1}^{*}$ is also an optimal solution for (17) and their optimal values are the same, we have
$x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}+\lambda_{w}\|Dx_{2}^{*}\|_{p}=x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}\ \Rightarrow\ \|Dx_{2}^{*}\|_{p}=0.$
Thus, $(x_{1}^{*},x_{2}^{*})$ is an optimal solution for (17). This completes the proof.

### v2

@@ 7 Conclusion
In this paper, we have analyzed the role of regularizations in direct and indirect data-driven control via a bi-level optimization framework. We prove that, after dropping some inner constraints, the bi-level optimization problem can be transformed into a single-level convex optimization problem. Moreover, regularizers developed from those single-level problems are exact penalty functions under certain conditions. Thus, using regularizers is a further convex relaxation with respect to the corresponding bi-level problem (after dropping constraints). Moreover, we have proposed a new variant, called A-DDPC, which uses an iterative algorithm to obtain a new data-driven control approach. Numerical simulations have demonstrated the superior performance of A-DDPC (38) with the more structured predictor.
Interesting future directions include: 1) investigating the closed-loop performance of the DDPC schemes with various regularizers, 2) characterizing the relationship between the performance of different DDPC variants and the degree of system nonlinearity, and 3) developing online trajectory library updated algorithms to ensure control performance for nonlinear systems, while preserving real-time efficiency. Another interesting direction for future work is to systematically decide the model order in (34a) for the proposed A-DDPC and validate it through real-world experiments, following settings similar to those in [elokda2021data, fawcett2022distributed, wang2023implementation, Shang2025Decentralized, lian2023adaptive, huang2021decentralized].
@@ Appendix A Exact penalty for quadratic optimization
A.1 A technical lemma
We here introduce Hoffman’s error bound for a system of linear equalities.
Theorem 5 (Hoffman’s error bound [hoffman1952approximate]).
Let the system of linear equalities $Ax=\bar{b}$ with $x\in\mathbb{R}^{n_{x}}$ be consistent and $\bar{\mathcal{F}}$ be its feasible solution set. Fix any $p\in[1,\infty]$. Then, there exists a constant $\bar{L}$, such that for any $x\in\mathbb{R}^{n}$ there exists a point $x^{*}\in\bar{\mathcal{F}}$ satisfying $\|x-x^{*}\|_{p}\leq\bar{L}\|Ax-\bar{b}\|_{p}.$
This theorem shows that if a vector does not satisfy the linear equality constraints, then its distance to the feasible set can be bounded by the size of the constraint violation. We now state a key technical lemma that will be used in the proof of Theorem 1.
Lemma 2 (Linear perturbation).
Let $A_{1}\in\mathbb{R}^{m\times n_{1}}$, $A_{2}\in\mathbb{R}^{m\times n_{2}}$, $D\in\mathbb{R}^{s\times n_{2}}$, and $b\in\mathbb{R}^{m}$. Suppose the set
$\mathcal{F}=\{(x_{1},x_{2})\mid A_{1}x_{1}+A_{2}x_{2}=b,\ \ Dx_{2}=0\}$
is non-empty. Fix any $p\in[1,\infty]$. Then there exists a finite constant $L>0$ such that for any pair $(x_{1},x_{2})$ satisfying $A_{1}x_{1}+A_{2}x_{2}=b$, there exists a pair $(\tilde{x}_{1},\tilde{x}_{2})$ satisfying $(\tilde{x}_{1},\tilde{x}_{2})\in\mathcal{F},\,\|x_{1}-\tilde{x}_{1}\|_{p}\leq L\|Dx_{2}\|_{p}.$
@@ A.1 A technical lemma
We can construct $A:=\begin{bmatrix}A_{1}&A_{2}\\ 0&D\end{bmatrix},\bar{b}:=\begin{bmatrix}b\\ 0\end{bmatrix}$ and then $\mathcal{F}$ is the set of feasible solutions for constraint $Ax=\bar{b}$. From Theorem 5, there exists $L$ such that, for any $x:=\textnormal{col}(x_{1},x_{2})$ there exists $\tilde{x}:=(\tilde{x}_{1},\tilde{x}_{2})\in\mathcal{F}$ with
$\|x_{1}-\tilde{x}_{1}\|_{p}$ $\leq\|x-\tilde{x}\|_{p}$
$\leq L\|\begin{bmatrix}A_{1}x_{1}+A_{2}x_{2}-b\\ Dx_{2}\end{bmatrix}\|_{p}=L\|Dx_{2}\|_{p},$
where the last equality comes from $A_{1}x_{1}+A_{2}x_{2}=b$. This completes the proof.
A.2 Proof of Theorem 1
Our proof below adapts the idea of partial calmness [ye1997exact] into the quadratic optimization problems in Theorem 1.
We let $f(x)\!:=\!x^{\top}Mx$ for simplicity. Since $\|\!\cdot\!\|_{p}$ is a norm, the constraint $\|Dx_{2}\|_{p}\!=\!0$ is equivalent to $Dx_{2}\!=\!0$. It suffices to prove that there exists a finite $\lambda_{w}^{*}>0$ such that for every $\lambda_{w}>\lambda_{w}^{*}$, every global minimizer $(x_{1},x_{2})$ of (18) must satisfy $Dx_{2}=0$. Indeed, if this holds, then all minimizers of (18) are feasible for (17). On the set $\{(x_{1},x_{2}):Dx_{2}=0\}$, the penalty term vanishes, so both problems have the same objective values and constraints. Therefore, the two problems have the same optimal value and the same set of optimal solutions.
Step 1: Every minimizer $(x_{1},x_{2})$ of (18) must satisfy $Dx_{2}=0$. We claim that there exist constants $\delta>0$ and $\lambda_{w}^{*}>0$ such that for any $\lambda_{w}>\lambda_{w}^{*}$, the inequality
$f(x_{1})+\lambda_{w}\|Dx_{2}\|_{p}\ \geq\ f(x_{1}^{*})+(\lambda_{w}-\lambda_{w}^{*})\|Dx_{2}\|_{p}$ (40)
holds for all feasible $(x_{1},x_{2})$ of (18) satisfying
$\|x_{1}-x_{1}^{*}\|_{p}\leq\delta,\qquad\|Dx_{2}\|_{p}\leq\delta,$ (41)
where $(x_{1}^{*},x_{2}^{*})$ is the given optimal solution of (17). We note that, the existence of such feasible solution is guaranteed as we can trivially choose $x_{1}=x_{1}^{*}$ and $x_{2}=x_{2}^{*}$.
Assuming (40) for now, fix any $\lambda_{w}>\lambda_{w}^{*}$. Since $(x_{1}^{*},x_{2}^{*})$ is feasible for (18) and satisfies $Dx_{2}^{*}=0$, it attains objective value $f(x_{1}^{*})$ in (18). Moreover, (40) implies that any feasible point in the neighborhood (41) has objective value at least $f(x_{1}^{*})$, with strict inequality whenever $Dx_{2}\neq 0$. Hence $(x_{1}^{*},x_{2}^{*})$ is a local minimizer of (18). Since (18) is convex, every local minimizer is global, so $(x_{1}^{*},x_{2}^{*})$ is a global minimizer of (18).
Now let $(\bar{x}_{1},\bar{x}_{2})$ be any global minimizer of (18). If $D\bar{x}_{2}\neq 0$, consider the convex combination
$(x_{1,t},x_{2,t}):=(1-t)(x_{1}^{*},x_{2}^{*})+t(\bar{x}_{1},\bar{x}_{2}),\qquad t\in(0,1),$
which is feasible to (18) for all $t\in[0,1]$. Since the objective is convex and both endpoints are global minimizers, each $(x_{1,t},x_{2,t})$ is also a global minimizer. Choose $t>0$ small enough so that (41) holds, i.e., $t\|\bar{x}_{1}-x_{1}^{*}\|_{p}\leq\delta,\,t\|D\bar{x}_{2}\|_{p}\leq\delta.$ Then $Dx_{2,t}=tD\bar{x}_{2}\neq 0$, and applying (40) gives $f(x_{1,t})+\lambda_{w}\|Dx_{2,t}\|_{p}\ \geq\ f(x_{1}^{*})+(\lambda_{w}-\lambda_{w}^{*})\|Dx_{2,t}\|_{p}\ >\ f(x_{1}^{*}),$ contradicting that $(x_{1,t},x_{2,t})$ is a global minimizer. Therefore, every global minimizer $(\bar{x}_{1},\bar{x}_{2})$ of (18) must satisfy $D\bar{x}_{2}=0$. This completes the proof once we establish (40).
Step 2: Proof of the key inequality (40). Since $x_{1}^{*}$ is an interior point of $\mathcal{X}$, there exists $r>0$ such that the $p$-norm ball $\mathcal{B}_{p}(x_{1}^{*};r)\subseteq\mathcal{X}$. Because $f$ is quadratic and $\mathcal{B}_{p}(x_{1}^{*};r)$ is bounded, $f$ is Lipschitz on this ball, i.e., there exists $\kappa_{r}>0$ such that
$|f(u)-f(v)|\leq\kappa_{r}\|u-v\|_{p},\quad\forall u,v\in\mathcal{B}_{p}(x_{1}^{*};r).$ (42)
Since $(x_{1}^{*},x_{2}^{*})$ is feasible for (17), the set $\mathcal{F}:=\{(x_{1},x_{2}):A_{1}x_{1}+A_{2}x_{2}=b,\ Dx_{2}=0\}$ is nonempty. Hence Lemma 2 yields a constant $L>0$ such that for any $(x_{1},x_{2})$ satisfying $A_{1}x_{1}+A_{2}x_{2}=b$, there exists $(\tilde{x}_{1},\tilde{x}_{2})\in\mathcal{F}$ with
$\|x_{1}-\tilde{x}_{1}\|_{p}\leq L\|Dx_{2}\|_{p}.$ (43)
Now take any feasible $(x_{1},x_{2})$ of (18) satisfying (41) and let $(\tilde{x}_{1},\tilde{x}_{2})\in\mathcal{F}$ be the corresponding perturbed point from Lemma 2. Then
$\|\tilde{x}_{1}-x_{1}^{*}\|_{p}\leq\|x_{1}-x_{1}^{*}\|_{p}+\|x_{1}-\tilde{x}_{1}\|_{p}\leq\delta+L\delta=r,$
so $\tilde{x}_{1}\in\mathcal{B}_{p}(x_{1}^{*};r)\subseteq\mathcal{X}$ via choosing $\delta:=r/(1+L)$. That illustrates $(\tilde{x}_{1},\tilde{x}_{2})$ is feasible for (17). By optimality of $(x_{1}^{*},x_{2}^{*})$ for (17),
$f(\tilde{x}_{1})\geq f(x_{1}^{*}).$
Moreover, since $x_{1}\in\mathcal{X}$, $\tilde{x}_{1}\in\mathcal{B}_{p}(x_{1}^{*};r)\subseteq\mathcal{X}$, and both are in $\mathcal{B}_{p}(x_{1}^{*};r)$ whenever (41) holds (note that $\delta<r$ from construction), the local Lipschitz bound (42) gives
$f(x_{1})\geq f(\tilde{x}_{1})-\kappa_{r}\|x_{1}-\tilde{x}_{1}\|_{p}\geq f(x_{1}^{*})-\kappa_{r}L\|Dx_{2}\|_{p},$
where we used (43) in the last step. Therefore, for any $\lambda_{w}\geq 0$,
$f(x_{1})+\lambda_{w}\|Dx_{2}\|_{p}\geq f(x_{1}^{*})+(\lambda_{w}-\kappa_{r}L)\|Dx_{2}\|_{p}.$
This is exactly (40) with $\lambda_{w}^{*}:=\kappa_{r}L$. This completes the proof.